# SCOPE: Stochastic Cartographic Occupancy Prediction Engine for Uncertainty-Aware Dynamic Navigation

### v1

@@ III-B Image-Based Prediction
An OGM can be considered a grayscale image, with the probability of occupancy defining the “color” in each cell of the regular grid. Using this interpretation, one could use image/video prediction algorithms to generate the next OGM given a short sequence of past OGMs. The prediction model (1) of image-based approaches is rewritten as
@@
$\mathbf{o}_{t-\tau:t}$ $=\ e(\mathbf{y}_{t-\tau:t}),$ (2b)
where $f_{\theta}(\cdot)$ is a deep neural network model, and $e(\cdot)$ is the conversion function to convert the lidar measurements to the binary OGMs in the robot’s local coordinate frame (i.e., using ray tracing). From this image-based model (2), we can easily see that image-based methods explicitly ignore the kinematics and dynamics of the robot and surrounding objects (i.e., they assume motion can be implicitly captured by very powerful network architectures and enough good data), and fail to provide a range of possible and reliable OGM predictions (i.e., they assume a deterministic future).
@@ III-C SCOPE++
Based on the limitations of image-based methods above, we argue that: 1) the future state of the environment explicitly depends on the motion of the robot itself, the motion of dynamic objects, and the state of static objects within the environment; 2) the future state of the environment is stochastic and unknown, and a range of possible future states helps provide robust predictions. With these two assumptions, we fully and explicitly exploit the kinematic and dynamic information of these three different types of objects, utilize the VAE-based network to provide stochastic predictions, and finally propose our novel stochastic OGM predictor SCOPE (shown in Fig. 2) to predict the future state of the environment. We also create a second variant called SCOPE++ that explicitly separates out static objects. Since SCOPE++ is a superset of SCOPE, the remainder of this section will focus on SCOPE++.
The prediction model (1) of our SCOPE++, outlined in Fig. 2, can be rewritten as
@@
$\mathbf{\hat{o}}_{t+1}=$ $\ h(\mathbf{o}_{t-\tau:t}),$ (3b)
$\mathbf{o}_{t-\tau:t}=$ $\ e({\mathbf{y}^{R}_{t-\tau:t}}),$ (3c)
$\mathbf{m}=$ $\ g({\mathbf{y}^{R}_{t-\tau:t}}),$ (3d)
$\mathbf{y}^{R}_{t-\tau:t}=$ $\ k(\mathbf{d}_{t-\tau:t}).$ (3e)
@@
In this framework, (3a) is the VAE predictor module from which we sample future maps $\mathbf{o}_{t+1}$. (3b) and (3c) represent the dynamic object module, where $h(\cdot)$ processes the time series data for dynamic objects, $e(\cdot)$ is the conversion function like (2b), and $R$ denotes the local coordinate frame of the robot at predicted time $t+n$. (3d) is the static objects module, where $g(\cdot)$ is the occupancy grid mapping function for static objects. Finally, (3e) is the robot motion module, where $k(\cdot)$ is the transformation function to compensate for robot motion.
@@
Note that (3) only predicts future states at the next time step $t+1$ and is used for training. To predict a multi-step future states at time $t+n$, we can easily utilize the autoregressive mechanism and feed the next-step prediction back $\mathbf{o}_{t+1}$ to our SCOPE/SCOPE++ network (3a) for $n-1$ time steps to predict the future states at time step $t+n$. Note that the prediction horizon $n$ could theoretically be any time step.
III-C1 Robot Motion Compensation $k(\cdot)$
To account for motion of the robot, we propose a simple and effective ego-motion compensation mechanism. We first predict the future pose of the robot at the end of our prediction horizon, time $t+n$. To do this, we use a constant velocity motion model, as it is the most widely used motion model for tracking [60] and often outperforms more state-of-the-art methods in general settings [61] when used on relatively short period (i.e., less than 1 s). Note that other more suitable robot motion models specific to particular robot models can be used to provide better robot pose predictions.
@@ III-C1 Robot Motion Compensation $k(\cdot)$
This results in a predicted pose for the robot $\mathbf{x}_{t+n}$, and we use this pose as the origin of our coordinate frame $R$ (where the x axis is forward and y axis is left, as is standard in mobile robotics [62]). We then transform all data into this frame, resulting in poses $\mathbf{x}_{t-\tau:t}^{R}$ and lidar scans $\mathbf{y}_{t-\tau:t}^{R}$. Using this data, we can generate the sequence of OGMS $\mathbf{o}_{t-\tau:t}^{R}$ using the conversion function $e(\cdot)$ by setting each lidar scan “hit” to 1 and all other cells to 0 (for more details, see [4]). The benefit of ego-motion compensation is that we can treat these lidar measurements from a moving lidar sensor as observations in common world-fixed frame $R$, which significantly reduces the difficulty of OGM predictions and improves accuracy. Note this only requires the robot to have accurate odometry over the history window $\tau$ (typically on the order of 1 s), but not accurate global localization.
III-C2 Dynamic Object Prediction $h(\cdot)$
@@
Tracking and predicting dynamic objects such as pedestrians is the hardest part of environmental prediction in complex dynamic scenes. It requires some techniques to process a set of time series data to capture the motion information. While the traditional particle-based methods [13] require explicitly tracking objects and treat each grid cell as an independent state, recent learning-based methods [14, 15, 16, 17, 18, 20, 21, 22, 25] prefer to use RNNs to directly process the observed time series OGMs. Based on these trends, we choose the most popular ConvLSTM unit to process the spatiotemporal OGM sequences ${\mathbf{o}_{t-\tau:t}}$. However, while other works [7, 15] explicitly decouple the dynamic and static/unknown objects and use different networks to process them separately, we argue that the motion of dynamic objects is related to their surroundings, and that explicit disentangling may lose some useful contextual information. For example, pedestrians walking through a narrow corridor are less likely to collide with or pass through surrounding walls. To exploit the useful contextual information between dynamic objects and their surrounding, we directly feed the observed OGMs $\mathbf{o}_{t-\tau:t}^{R}$ into a ConvLSTM unit $h(\cdot)$ and implicitly predict the future state ${\mathbf{\hat{o}}_{t+1}}$ of dynamic (and static) objects.
@@ III-C3 Static Object Segmentation $g(\cdot)$
While predicting dynamic objects plays a key role in environmental prediction, paying extra attention to static objects is also important to improve prediction accuracy. The main reason is that the area occupied by static objects is much larger than that of dynamic objects, as shown in Fig. 2, where static objects such as walls much larger than dynamic objects such as pedestrians, which are sparse and scattered point clusters. Another reason is that static objects maintain their shape and position over time, contributing to the scene geometry and giving a global view of the surroundings.
To account for static objects, we utilize a local static environment map $\mathbf{m}$ as a prediction for future static objects, which is a key contribution of our work. We generate this local environment map $\mathbf{m}$ using a GPU-accelerated implementation [^2] https://github.com/TempleRAIL/occupancy_grid_mapping_torch of the standard inverse sensor model [63] for $g(\cdot)$ that parallelizes independent cell state update operations by taking advantage of the additive update rule (for more details, see [4]). This Bayesian approach generates a robust estimate of the local map, where dynamic objects are treated as noise data and removed over time.
@@ III-C4 VAE Predictor
The dynamic and static prediction functions from (3b) and (3d) yield one estimate of the future environment. However, it is unlikely that this exact map will be correct. To account for this, we wish to generate a range of possible future environment states that capture the uncertainty of the environment. We do this using a variational autoencoder (VAE), a generative machine learning technique that creates samples that are “similar” to the predicted OGM $\mathbf{\hat{o}}_{t+1}$.
To represent the stochasticity of environment states, we assume that environment states $\mathbf{o}_{t-\tau:t+n}$ are generated by some unobserved, random, latent variable $\mathbf{z}$ that follow a prior distribution $p_{\theta}(\mathbf{z})$. Our VAE predictor model (3a) can be rewritten as
@@
Finally, we use mini-batching, Monte-Carlo estimation, and reparameterization tricks to calculate the stochastic gradients of the ELBO (5[64], and obtain the optimized model parameters $\phi$ and $\theta$. Using them, our VAE predictor can integrate the predicted features $\{\mathbf{\hat{o}}_{t+1},\mathbf{m}\}$ of dynamic and static objects, and output a probabilistic estimate of the future OGM states with uncertainty awareness.
@@ IV Software Optimization
We demonstrated in our previous work [4] (results also in Section V) that SCOPE and SCOPE++ are able to accurately predict maps. However, the use of multiple neural networks is resource intensive. For example, the VAE used to generate samples to provide uncertainty estimates is so memory-intensive that we can only generate 8 samples using an NVIDIA Jetson AVG Xavier embedded device [4]. [^3] Generating 8 samples is a trade-off in running speed and memory consumption when running with other navigation-related packages such as amcl, move_base, and other sensor driver ROS packages. In order to make our proposed SCOPE/SCOPE++ predictors work efficiently in different real-world robotics applications, we decide to optimize them on the software side.
Runtime 9.04% 53.41% 20.93% 16.62%
Memory Usage 0.32% 1.28% 48.38% 50.02%
@@ IV-A Resource Utilization
Table I details the running performance percentage of each module in the SCOPE++ predictor (generating 8 samples) in terms of runtime and memory usage on our embedded computer. We can easily see that the runtime is mainly limited by the static and dynamic object modules (with the ConvLSTM unit), while the memory usage is mainly limited by the dynamic object (with ConvLSTM unit) and VAE modules. This analysis explains why our proposed SCOPE predictor without static object module is much faster than our SCOPE++ predictor.
Since the key ConvLSTM unit in the dynamic object module is already an accelerated RNN unit (compared to the Vanilla RNN unit), it is difficult to further optimize this module. Therefore, we choose to perform software optimization on the static object module and VAE module to increase running speed and reduce memory consumption while maintaining accurate prediction performance and the stochastic nature of our SCOPE/SCOPE++ predictors.
According to Table I and our network architecture design, a straightforward method is to remove the static object module (i.e., use SCOPE instead of SCOPE++), especially since the two variants achieve comparable prediction accuracy. Next, we will look at replace the VAE network by utilizing knowledge distillation techniques to compress our SCOPE network to maintain its prediction performance, and then quantify the prediction uncertainty of the SCOPE to provide uncertainty estimates to maintain its stochastic nature.
@@ IV-B Knowledge Distillation
We use the SCOPE network instead of SCOPE++ network as our teacher network for knowledge distillation, shown in Fig. 2. There are two reasons why we choose the fast and less accurate SCOPE network rather than the complex and more accurate SCOPE++ network as the teacher network. First, as Cho et al. [44] suggested, larger or more accurate models generally do not make better teachers in the knowledge distillation, especially when the student’s ability is too low to successfully imitate the teacher. Second, Beyer et al. [66] suggested that keeping “consistent” teaching can improve knowledge distillation. Our SO-SCOPE student network follows the SCOPE network design, which is almost the same as the SCOPE network except that a single convolutional layer is used instead of the VAE network (i.e., the input view of the convolutional layer is consistent with the VAE network). Therefore, this structural similarity and consistency improves knowledge transfer from the teacher network to the student network. This will accelerate the running speed and also reduce the memory consumption.
To train our SO-SCOPE student network, we directly use the output of our trained SCOPE teacher model as “soft” labels instead of using ground truth labels, which is following the key idea of knowledge distillation. Note that since our SCOPE predictor is stochastic and generates different samples even with the same input, this stochastic sample output expands the original training dataset and further improves the generalization ability of the SO-SCOPE predictor. We use the binary cross entropy (BCE) loss function to train our SO-SCOPE student network and get our final SO-SCOPE predictor to provide deterministic OGM predictions. After completing the knowledge distillation training, we directly use the output of the SO-SCOPE predictor to provide OGM prediction information. However, this distilled network is now deterministic, i.e., each input will generate a single output. Our next step will address this limitation.
@@
Figure 3: Uncertainty quantification visualization of predicted OGM cell bins for different prediction timesteps. Grey histograms are the raw predicted OGM cell data of our SCOPE++ predictor. Blue curves are the connecting curves of these raw predicted OGM cell data. Red curves are the fitting curves from our proposed mixture model of truncated normal and skew-Cauchy distributions.
@@ IV-C Uncertainty Quantification
To modify SO-SCOPE, we will characterize the statistics of the VAE from our trained SCOPE++ predictor and then combine this data with the output of distillation network (which should generate an OGM that is similar to an output of the VAE) to draw samples with identical statistics to our original network but in a much more memory- and time-efficient manner. Note that using SOGMP++ here is a reasonable choice because our SCOPE++ predictor provides a more accurate and comprehensive uncertainty estimate than SCOPE. To characterize the VAE, we collected 8000 different input sequences to the ConvLSTM from the OGM-Turtlebot2 test dataset (see Section V-A) and generated 32 VAE samples for each input, giving us a total of $256,000$ output samples for each prediction horizon $T=1,\ldots,\tau$.
@@
Our analysis relies on two factors. First, in an OGM, the probability of occupancy for each cell is independent of all other cells [63]. In other words, for two cells $c$ and $c^{\prime}$, then $p(c\textrm{ is occupied and }c^{\prime}\textrm{ is occupied})=p(c\textrm{ is occupied})p(c^{\prime}\textrm{ is occupied})$. Second, based on our use of a VAE, we make the assumption that the distribution of occupancy probabilities in the VAE output depends on the occupancy probability of the input. This is reasonable since a VAE is trained to have similar outputs be near to one another in the learned latent space. Let $\hat{c}_{t+T}=p(c\textrm{ is occupied at time }t+T)$ denote the probability of occupancy of cell $c$ predicted by the ConvLSTM layer $T$ steps into the future (i.e., we use the average of the occupancy probabilities predicted by the VAE for the same cell as its estimate) and $\tilde{c}_{t+T}$ be the probability of occupancy predicted for that same cell by the VAE. Then we assume $p(\tilde{c}_{t+T}\mid\hat{c}_{t+T})\neq p(\tilde{c}_{t+T})$.
@@
Based on these factors, we will examine trends in the VAE probabilities $\tilde{c}_{t+T}$ (for $T=1,\ldots,\tau$) conditioned on the value of $\hat{c}_{t+T}$. We use 15 bins for $\hat{c}_{t+T}$, evenly spaced from 0 to 1 (since the probability range of OGM cell is in $[0,1]$). With these binned OGM cell samples, we visualize the distribution of $\tilde{c}_{t+T}$ for each $\hat{c}_{t+T}$ bin, as shown in Fig. 3. We can see a few useful trends in the data. First, cells that are predicted to be empty by the ConvLSTM (i.e., in bin 0) remain empty for nearly all VAE samples. This trend holds during the full prediction window, with only a slight increase in the probability of occupancy at $T$ gets larger. Second, we see that the peak of the distribution for $\tilde{c}_{t+T}$ remains very close to the value of $\hat{c}_{t+1}$. This implies that the VAE generates a consistent distribution around the ConvLSTM output. Third, we see heavier tails in the distributions as $T$ increases. This makes sense as a longer horizon into the future leads to further uncertainty and further deviation from the current time. Fourth, we see the distributions skew towards 0 as the prediction horizon $T$ increases. This makes sense as cells are more likely to become empty as people move about in the scene. Fifth, for situations where $\hat{c}_{t+1}$ is low, we see a peak in the distribution of $\tilde{c}_{t+T}$ at 0. This is cause by uncertain objects moving out of a cell, leaving it empty.
@@
Our goal is to fit a model to the resulting histograms. Our model must be able to capture the heavy tails, skew, and the peak near $\tilde{c}=0$. Based on this, we use a mixture model of a truncated normal and a skew-Cauchy distribution to describe the distribution of $\tilde{c}_{t+T}$ values, which can be expressed mathematically as:

### v2

@@ III-B Image-Based Prediction
An OGM can be considered a grayscale image, with the probability of occupancy defining the “color” in each cell of a regular grid. Using this interpretation, image/video prediction algorithms could be used to generate the next OGM given a short sequence of past OGMs [17, 19, 20]. The prediction model (1) of image-based approaches is rewritten as
@@
$\mathbf{o}_{t-\tau:t}$ $=\ \psi(\mathbf{y}_{t-\tau:t}),$ (2b)
where $f_{\theta}(\cdot)$ is a DNN model, and $\psi(\cdot)$ is the conversion function to convert the lidar measurements to the binary OGMs in the robot’s local frame (i.e., using ray tracing). From this image-based model (2), we can see that image-based methods explicitly ignore the kinematics and dynamics of the robot and surrounding objects (i.e., they assume motion can be implicitly captured by powerful network architectures and enough good data), and fail to provide a range of possible and reliable OGM predictions (i.e., they assume a deterministic future).
@@ III-C SCOPE++
Based on the limitations of image-based methods above, we argue that: 1) the future state of the environment explicitly depends on the motion of the robot itself, the motion of dynamic objects, and the state of static objects within the environment; 2) the future state of the environment is stochastic and unknown, and a range of possible future states helps provide robust predictions. With these two assumptions, we fully and explicitly exploit the kinematic and dynamic information of these three different types of objects, utilize the VAE-based network to provide stochastic predictions, and finally propose our novel stochastic OGM predictors SCOPE series (shown in Fig. 2) to predict the future state of the environment. We first create a variant called SCOPE++ that explicitly separates out static objects. Since SCOPE++ is a superset of SCOPE, the rest of this section will focus on SCOPE++. The prediction model (1) of our SCOPE++, outlined in Fig. 2, can be rewritten as
@@
$\mathbf{\hat{o}}_{t+1}=\ \kappa(\mathbf{o}_{t-\tau:t}),$ $\ \mathbf{o}_{t-\tau:t}=\ \psi({\mathbf{y}^{R}_{t-\tau:t}}),$ (3b)
$\mathbf{m}=$ $\ g({\mathbf{y}^{R}_{t-\tau:t}}),$ (3c)
$\mathbf{y}^{R}_{t-\tau:t}=$ $\ \Lambda(\mathbf{d}_{t-\tau:t}).$ (3d)
@@
In this framework, (3a) is the VAE predictor module from which we sample future maps $\mathbf{o}_{t+1}$. (3b) represent the dynamic object module, where $\kappa(\cdot)$ processes the time series data for dynamic objects, $\psi(\cdot)$ is the conversion function like (2b), and $R$ denotes the local coordinate frame of the robot at predicted time $t+n$. (3c) is the static objects module, where $g(\cdot)$ is the occupancy grid mapping function for static objects. Finally, (3d) is the robot motion module, where $\Lambda(\cdot)$ is the transformation function to compensate for robot motion.
@@
Note that (3) only predicts future states at the next time step $t+1$ and is used for training. To predict a multi-step future state at time $t+n$, we can easily utilize the autoregressive mechanism and feed the next-step prediction back $\mathbf{o}_{t+1}$ to our SCOPE/SCOPE++ network (3a) for $n-1$ time steps to predict the future states at time step $t+n$. Note that the prediction horizon $n$ could theoretically be any time step.
III-C1 Robot Motion Compensation $\Lambda(\cdot)$
@@
To account for robot motion, we propose a simple and effective ego-motion compensation mechanism. We first predict the future pose of the robot at the end of our prediction horizon, time $t+n$. To do this, we use a constant velocity motion model, as it is the most widely used motion model for tracking [62] and often outperforms more state-of-the-art methods in general settings [63] when used on a relatively short period (i.e., less than 1 s). Note that other more appropriate robot motion models tailored to specific robot models can be used to provide better ego-motion compensation. This results in a predicted pose for the robot $\mathbf{x}_{t+n}$, and we use this pose as the origin of our coordinate frame $R$ (where the x-axis is forward and the y-axis is left, as is standard in mobile robotics [64]). We then transform all data into this frame, resulting in poses $\mathbf{x}_{t-\tau:t}^{R}$ and lidar scans $\mathbf{y}_{t-\tau:t}^{R}$. Using this data, we can generate the sequence of OGMS $\mathbf{o}_{t-\tau:t}^{R}$ using the conversion function $\psi(\cdot)$ by setting each lidar scan “hit” to 1 and all other cells to 0 (for more details, see [4]). The benefit of ego-motion compensation is that we can treat these lidar measurements from a moving lidar sensor as observations in common world-fixed frame $R$, which significantly reduces the difficulty of OGM predictions and improves accuracy. Note this only requires the robot to have accurate odometry over the history window $\tau$ (typically on the order of 1 s), but not accurate global localization.
III-C2 Dynamic Object Prediction $\kappa(\cdot)$
@@
Tracking and predicting dynamic objects such as pedestrians is the hardest part of environmental prediction in complex dynamic scenes. It requires some techniques to process a set of time series data to capture the motion information. While the traditional particle-based methods [16] require explicitly tracking objects and treat each grid cell as an independent state, recent learning-based methods [17, 18, 19, 20, 8, 22, 23, 24, 27] prefer to use RNNs to directly process the observed time series OGMs. Based on these trends, we choose the most popular ConvLSTM unit to process the spatiotemporal OGM sequences ${\mathbf{o}_{t-\tau:t}}$. However, while other works [7, 18] explicitly decouple the dynamic and static/unknown objects and use different networks to process them separately, we argue that the motion of dynamic objects is related to their surroundings, and that explicit disentangling may lose some useful contextual information. For example, pedestrians walking through a narrow corridor are less likely to collide with or pass through surrounding walls. To exploit the useful contextual information between dynamic objects and their surrounding, we directly feed the observed OGMs $\mathbf{o}_{t-\tau:t}^{R}$ into a ConvLSTM unit $\kappa(\cdot)$ and implicitly predict the future state ${\mathbf{\hat{o}}_{t+1}}$ of dynamic (and static) objects. [^2] Note that since the VAE is designed to provide uncertainty estimates of future states (while the static object module is designed to provide static information), our supervised learning mechanism will implicitly regularize the output of the ConvLSTM unit to contain the future states of dynamic objects.
Runtime 9.04% 53.41% 20.93% 16.62%
Memory Usage 0.32% 1.28% 48.38% 50.02%
@@ III-C3 Static Object Segmentation $g(\cdot)$
While predicting dynamic objects plays a key role in environmental prediction, paying extra attention to static objects is also important to improve prediction accuracy. The main reason is that the area occupied by static objects is much larger than that of dynamic objects, as shown in Fig. 2, where static objects such as walls are much larger than dynamic objects such as pedestrians, which are sparse and scattered point clusters. Another reason is that static objects maintain their shape and position over time, contributing to the scene geometry and giving a global view of the surroundings. To account for static objects, we utilize a local static environment map $\mathbf{m}$ as a prediction for future static objects, which is a key contribution of our work. We generate this local environment map $\mathbf{m}$ using a GPU-accelerated implementation [^3] https://github.com/TempleRAIL/occupancy_grid_mapping_torch of the standard inverse sensor model [65] for $g(\cdot)$ that parallelizes independent cell state update operations by taking advantage of the additive update rule (for more details, see [4]). This Bayesian approach generates a robust estimate of the local map, where dynamic objects are treated as noise data and removed over time.
@@ III-C4 VAE Predictor
The dynamic and static prediction functions from (3b) and (3c) yield one estimate of the future environment. However, it is unlikely that this exact map will be correct. To account for this, we wish to generate a range of possible future environment states that capture the uncertainty of the environment. We do this using a variational autoencoder (VAE), a generative machine learning technique that creates samples that are “similar” to the predicted OGM $\mathbf{\hat{o}}_{t+1}$.
To represent the stochasticity of environment states, we assume that environment states $\mathbf{o}_{t-\tau:t+n}$ are generated by some unobserved, random, latent variable $\mathbf{z}$ that follow a prior distribution $p_{\theta}(\mathbf{z})$. Our VAE model (3a) can be rewritten as
@@
We use mini-batching, Monte-Carlo estimation, and reparameterization tricks to calculate the gradients of the ELBO (5[66], and obtain the optimized model parameters $\phi$ and $\theta$. Finally, our VAE can integrate the predicted features $\{\mathbf{\hat{o}}_{t+1},\mathbf{m}\}$ of dynamic and static objects, and output a probabilistic estimate of future OGM states with uncertainty awareness.
@@ IV Software Optimization
We demonstrated in our previous work [4] (results also in Section V) that SCOPE and SCOPE++ can accurately predict OGMs. However, the use of multiple neural networks is resource-intensive. For example, the VAE used to generate samples to provide uncertainty estimates is so memory-intensive that we can only generate 8 samples using an NVIDIA Jetson AVG Xavier embedded device [4]. [^4] Generating 8 samples is a trade-off in running speed and memory consumption when running with other navigation-related packages such as amcl, move_base, and other sensor driver ROS packages. To make our proposed SCOPE/SCOPE++ work efficiently in resource-constrained robots, we decide to optimize them on the software side.
@@ IV-A Resource Utilization
Table I details the running performance percentage of each module in the SCOPE++ predictor (generating 8 samples) in terms of runtime and memory usage on our embedded computer. We can easily see that the runtime is mainly limited by the static and dynamic object modules (with the ConvLSTM unit), while the memory usage is mainly limited by the dynamic object (with ConvLSTM unit) and VAE modules. This analysis explains why our proposed SCOPE predictor without the static object module is much faster than our SCOPE++ predictor.
Since the key ConvLSTM unit in the dynamic object module is already an accelerated RNN unit (compared to the Vanilla RNN unit), it is difficult to further optimize this module. Therefore, we choose to perform software optimization on the static object module and VAE module to increase running speed and reduce memory consumption while maintaining the accurate prediction and stochasticity of our SCOPE/SCOPE++.
According to Table I and our network architecture design, a straightforward method is to remove the static object module (i.e., use SCOPE instead of SCOPE++), especially since the two variants achieve comparable prediction accuracy. Next, we will look at replacing the VAE network by utilizing knowledge distillation techniques to compress our SCOPE network to maintain its prediction performance, and then quantify the prediction uncertainty of the SCOPE to provide uncertainty estimates to maintain its stochastic nature.
@@ IV-B Knowledge Distillation
We use the SCOPE instead of the SCOPE++ as our teacher network for knowledge distillation, shown in Fig. 2. There are two reasons why we choose the faster and less accurate SCOPE network rather than the complex and more accurate SCOPE++ network as the teacher network. First, as Cho et al. [46] suggested, larger or more accurate models generally do not make better teachers in the knowledge distillation, especially when the student’s ability is too low to successfully imitate the teacher. Second, Beyer et al. [68] suggested that keeping “consistent” teaching can improve knowledge distillation. Our SO-SCOPE student network follows the SCOPE network design, which is almost the same as the SCOPE network except that a single convolutional layer is used instead of the VAE network (i.e., the input view of the convolutional layer is consistent with the VAE network). Therefore, this structural similarity and consistency improves knowledge transfer from the teacher network to the student network. This will accelerate the running speed and also reduce the memory consumption.
To train our SO-SCOPE student network, we directly use the output of our trained SCOPE teacher model in the OGM-Turtlebot2 training dataset (see Section V-A) as “soft” labels instead of using ground truth labels, which follow the key idea of knowledge distillation. To speed up training, we use 1 random sample output from SCOPE as a “soft” label in each training iteration. Note that since our SCOPE is stochastic and generates different samples even with the same input, this stochastic sample output expands the original training dataset (i.e., data augmentation) and further improves the generalization ability of the SO-SCOPE. We use the binary cross entropy (BCE) loss function to train our SO-SCOPE student network and get our final SO-SCOPE predictor to provide deterministic OGM predictions. After completing the knowledge distillation training, we directly use the output of the SO-SCOPE predictor to provide OGM prediction information. However, this distilled network is now deterministic, i.e., each input will generate a single output. Our next step will address this limitation.
@@
Figure 3: Uncertainty quantification of the predicted OGM cell bins (each column is one occupancy bin $\tilde{c}_{t+1}$) at different prediction time steps (each row is one timestep $T$). Grey histograms are the raw predicted OGM cell data of our SCOPE++ predictor. Blue curves are the connecting curves of these raw predicted OGM cell data. Red curves are the fitting curves from our proposed mixture model of truncated normal and skew-Cauchy distributions in (6).
@@ IV-C Uncertainty Quantification
To modify SO-SCOPE, we will characterize the statistics of the VAE from our trained SCOPE++ predictor and then combine this data with the output of the distillation network (which should generate an OGM that is similar to an output of the VAE) to draw samples with identical statistics to our original network but in a much more memory- and time-efficient manner. Note that using SCOPE++ here is a reasonable choice because our SCOPE++ predictor provides a more accurate and comprehensive uncertainty estimate than SCOPE. To characterize the VAE, we collected $8,000$ different input sequences to the ConvLSTM from the OGM-Turtlebot2 testing dataset (see Section V-A) and generated 32 VAE samples for each input, giving us a total of $256,000$ output samples for each prediction horizon $T=1,\ldots,\tau$. Although using the OGM-Turtlebot2 testing dataset may raise the issue of data leakage, we used the testing dataset to characterize the VAE because: 1) SCOPE++ is well trained on the training dataset and, as a result, the uncertainty of SCOPE++’s output in the training dataset is very small and the uncertainty statistics in the training dataset do not reflect the true uncertainty statistics of the SCOPE++; in contrast, the uncertainty statistics in the testing dataset reflect the true uncertainty statistics of the SCOPE++ very well; and 2) the key aim of proposing our software-optimized SO-SCOPE predictor is to optimize and accelerate the trained resource-intensive SCOPE/SCOPE++ predictors to allow them to run in real time on resource-constrained robots. It is reasonable to choose a testing dataset to characterize uncertainty statistics because we already have the predictors that need to be accelerated and we can easily use them to collect real data statistics from the application environment.
@@
Our analysis relies on two factors. First, in an OGM, the probability of occupancy for each cell is independent of all other cells [65]. In other words, for two cells $c$ and $c^{\prime}$, then $p(c\textrm{ is occupied and }c^{\prime}\textrm{ is occupied})=p(c\textrm{ is occupied})p(c^{\prime}\textrm{ is occupied})$. Second, based on our use of a VAE, we make the assumption that the distribution of occupancy probabilities in the VAE output depends on the occupancy probability of the input. This is reasonable since a VAE is trained to have similar outputs near to one another in the learned latent space. Let $\hat{c}_{t+T}=p(c\textrm{ is occupied at time }t+T)$ denote the probability of occupancy of cell $c$ predicted by the ConvLSTM layer $T$ steps into the future (i.e., we use the average of the occupancy probabilities predicted by the VAE for the same cell as its estimate) and $\tilde{c}_{t+T}$ be the probability of occupancy predicted for that same cell by the VAE. Then we assume $p(\tilde{c}_{t+T}\mid\hat{c}_{t+T})\neq p(\tilde{c}_{t+T})$.
@@
Based on these factors, we will examine trends in the VAE probabilities $\tilde{c}_{t+T}$ (for $T=1,\ldots,\tau$) conditioned on the value of $\hat{c}_{t+T}$. We use 15 bins for $\hat{c}_{t+T}$, evenly spaced from 0 to 1 (since the probability range of OGM cell is in $[0,1]$). With these binned OGM cell samples, we visualize the distribution of $\tilde{c}_{t+T}$ for each $\hat{c}_{t+T}$ bin, as shown in Fig. 3. We can see a few useful trends in the data. First, cells that are predicted to be empty by the ConvLSTM (i.e., in bin 0) remain empty for nearly all VAE samples. This trend holds during the full prediction window, with the occupancy probability increasing only slightly as $T$ becomes larger. Second, the peak of the distribution for $\tilde{c}_{t+T}$ remains very close to the value of $\hat{c}_{t+1}$. This implies that the VAE generates a consistent distribution around the ConvLSTM output. Third, heavier tails in the distributions as $T$ increases. This makes sense as a longer horizon into the future leads to further uncertainty and further deviation from the current time. Fourth, the distributions skew towards 0 as the prediction horizon $T$ increases. This makes sense as cells are more likely to become empty as people move about in the scene. Fifth, for situations where $\hat{c}_{t+1}$ is low, we see a peak in the distribution of $\tilde{c}_{t+T}$ at 0. This is caused by uncertain objects moving out of a cell, leaving it empty.
@@
Our goal is to fit a model to the resulting histograms. Our model must be able to capture the heavy tails, skew, and the peak near $\tilde{c}=0$. Based on this, we use a mixture model of a truncated normal and a skew-Cauchy distribution to describe the distribution of $\tilde{c}_{t+T}$ values, which can be expressed as: